# A Hybrid Strategy for Aggregated Probabilistic Forecasting and Energy Trading in HEFTCom2024

### v1

@@ Abstract
Obtaining accurate probabilistic energy forecasts and making effective decisions amid diverse uncertainties are routine challenges in future energy systems. This paper presents the solution of team GEB, which ranked 3rd in trading, 4th in forecasting, and 1st among student teams in the IEEE Hybrid Energy Forecasting and Trading Competition 2024 (HEFTCom2024). The solution provides accurate probabilistic forecasts for a wind-solar hybrid system, and achieves substantial trading revenue in the day-ahead electricity market. Key components include: (1) a stacking-based approach combining sister forecasts from various Numerical Weather Predictions (NWPs) to provide wind power forecasts, (2) an online solar post-processing model to address the distribution shift in the online test set caused by increased solar capacity, (3) a probabilistic aggregation method for accurate quantile forecasts of hybrid generation, and (4) a stochastic trading strategy to maximize expected trading revenue considering uncertainties in electricity prices. This paper also explores the potential of end-to-end learning to further enhance the trading revenue by adjusting the distribution of forecast errors. Detailed case studies are provided to validate the effectiveness of these proposed methods. Code for all mentioned methods is available for reproduction and further research in both industry and academia.
@@ 1 Introduction
Accurately forecasting renewable energy production and enabling timely decision-making are crucial for the transition to resilient, low-carbon power systems [1, 2, 3]. Despite continuous advancements in forecasting and decision-making techniques are reported in academia [4], significant challenges persist in addressing the uncertainties and complexities inherent in practical engineering applications. To bridge the gap between academic research and industrial practice, the Hybrid Energy Forecasting and Trading Competition 2024 (HEFTCom2024) challenges participants to provide quantile forecasts for a hybrid power plant which consists of the Hornsea 1 offshore wind farm and solar capacity in East England, and to trade the energy in a day-ahead electricity market.
@@
Our performance can be attributed to two key factors. On one hand, the primary wind and solar power forecasts are provided by Gradient Boosting Decision Trees (GBDTs), which is a popular ensemble learning approach that has demonstrated strong performance in previous energy forecasting competitions, such as Global Energy Forecasting Competitions (GEFCom) 2017 and GEFCom2014 [5, 6, 7, 8, 9]. We also incorporated several practical techniques including feature engineering with spatial and temporal meteorological information, hyperparameter optimization and ensemble learning. These techniques are widely recognized as effective strategies in forecasting practice [10, 11]. On the other hand, building on these proven techniques, we designed series tailored approaches to address the specific challenges posed by HEFTCom2024. To clarify the motivation behind our approaches, the following introduce these challenges or characteristics of HEFTCom2024, along with our corresponding solutions.
@@
(1) Multiple NWP sources. To reduce the forecast variance, we developed a stacking-based ensemble method to combine sister forecasts trained on two different NWP sources provided by HEFTCom2024. Stacking is a widely used ensemble learning technique that improves forecasting performance by leveraging the complementary strengths of diverse base models [12]. Stacking typically involves integrating models with different structures, such as combining Support Vector Regression (SVR) and Random Forest for day-ahead wind power forecasting [13], or combining Long Short Term Memory (LSTM) and Informer networks for solar power forecasting [14]. In contrast to heterogeneous ensembles, combining sister forecasts is also demonstrated effective in power load forecasting tasks [15], where sister forecasts refer to predictions generated from a series of models that share similar architectures but are trained using different input sources or variable selections [16]. In our solution, we stack sister forecasts trained on different NWP sources to exploit their spatial complementarity, as we observed discrepancies in grid point density and coverage area in the NWP data provided for the Hornsea 1 region, as shown in Section 2. Rather than directly merging multiple NWP sources into a single forecasting model, stacking sister models offers an additional advantage—it mitigates the risk of missing data from one NWP source during online testing. Since the two NWPs cover different spatial regions, embedding both into a single model would make the model vulnerable to data gaps if one NWP source becomes unavailable. In contrast, with a stacked ensemble, one sister model can still generate forecasts even when the other fails.
@@
(2) Solar distribution shift. The solar capacity in East England was observed from 2,609 MWp to 2,741 MWp as of February 19, 2024. Since the offline dataset only reflects capacities up to 2,609 MWp, the forecasting models solely trained on offline dataset become less effective at online test-set. Distribution shift is a typical challenge in forecasting practice. Common methods to address distribution shifts include oversampling, transfer learning, and post-processing. Oversampling replicates limited new samples to construct an augmented training set, which can alleviate overfitting to outdated data but may disrupt temporal continuity and lead to overfitting on the limited new samples or being overwhelmed by historical samples [17, 18]. Transfer learning aim to align the model outputs with the new data distribution through fine-tuning [19] or domain adaptation [20], but the structure of decision trees does not support typical transfer learning paradigms. Post-processing operates directly on model outputs, which is effective for linear shifts and does not require retraining the model. While they assume the existence of a known and simple shift pattern, such assumptions are valid in our case: the increase in solar capacity is both known and induces a near-linear expansion in power generation. We combined online learning with post-processing to dynamically fit the new distribution. Given the limited online samples, we adopted a polynomial post-processing model trained using LASSO-penalized quantile regression [21] for calibration, which prevents overfitting of higher-order terms by penalizing the model complexity. While wind power forecasts may also be subject to distributional changes, such shifts tend to be less systematic and harder to characterize without prior knowledge. Applying nonlinear corrections to wind power forecasts under sparse online data conditions will introduce additional risks of overfitting. Consequently, we restrict the use of online post-processing to solar power forecasts in the final solution.
@@
(3) Requirement to provide quantile forecasts of total generation. Directly predicting the distribution of total generation would hinder targeted post-processing, as Hornsea 1 encountered planned outage that were not present in the offline dataset, and the distribution of solar power generation also exhibited a shift. Therefore, forecasting wind and solar power separately is necessary and also help reduces the capacity of the forecasting model. While summing the forecast results quantile by quantile offers simplicity, it theoretically sacrifices the accuracy. Discrete convolution can be used to combine multiple cumulative distribution functions (CDFs) to obtain the distribution of the sum [22]. Copula-based discrete convolution methods are further researched to aggregate the uncertainties of wind power [23] and load demand [24] in power system. Inspired by these theoretical frameworks and applications, we developed an aggregated probabilistic forecasting approach by first predicting the CDF of wind and solar power separately and then reconstructing the distribution of total generation via discrete convolution. The CDF is estimated by learning the dense quantiles. This approach allows for an accurate estimation of the overall generation distribution in a parameter-free manner. Furthermore, this approach benefits from powerful models like LightGBM [25] or CatBoost [26], which directly support learning the pinball loss.
@@
(4) Requirement to trade energy in the day-ahead electricity market. The trading revenue is influenced by the uncertainties in day-ahead prices, imbalance prices, and power generation of the hybrid power plant. Due to limited information about the market and unknown market imbalances, price forecasting is challenging. However, statistically, the mean of the price spread follows a seasonal pattern throughout the day. To utilize this regularity, we developed a stochastic trading method based on uncertain price spread and deterministic power forecasts to maximize the expected trading revenue. We derived a revenue loss from the inaccuracies in both power forecasts and price spread forecasts to guide the training of the deterministic power forecasting model. This approach provides a simple and robust strategy that avoids the need for training price forecasting models. Additionally, we propose a further enhancement to improve the trading revenue through end-to-end learning (E2EL). E2EL refers to designing forecasting models with a focus on downstream decision-making tasks rather than solely minimizing statistical errors. E2EL is explored in related literature for unit commitment [27], economic dispatch [28], and energy arbitrage [29] tasks. The key to achieving superior decision-making through E2EL lies in leveraging the asymmetric impact of forecast errors on downstream benefits, which is overlooked in accuracy-oriented forecasting. We also discovered this characteristic of error asymmetry in the market rules set by HEFTCom2024, and case studies demonstrate that E2EL can enhance trading revenue by adjusting the distribution preferences of forecast errors.
@@ 2.3 Provided Data
HEFTCom2024 provides two meteorological data sources for forecasting: Deutscher Wetterdienst (DWD) and the Global Forecast System (GFS). The grid coordinates of the historical meteorological data for Hornsea 1 and East England photovoltaic (PV) plants are shown in Figure 1. For Hornsea 1, 36 coordinates are provided from the DWD, and 9 coordinates are provided from the GFS. For the solar power plants, 20 coordinates are provided from both DWD and GFS sources. The data types include wind speed and wind direction at 100m and 10m heights, temperature, solar radiation, cloud cover, and other meteorological data, with hourly resolution. The data are corrected every six hours, resulting in multiple corrected data points for the same valid time. The historical power data for Hornsea 1 and East England PV plants are given at half-hour intervals [30]. During the online testing phase, the wind power data for Hornsea 1 is typically updated with a one-week lag, while the solar power plant data is updated with a one-day lag. HEFTCom2024 also provides day-ahead and imbalance price data in Great Britain’s wholesale electricity market for the trading track. The aforementioned data span from September 20, 2020, to January 18, 2024. The evaluation period for the competition is from February 20, 2024, to May 19, 2024.
@@ 3.1 Overview
The overall framework of the developed methodology is illustrated in Figure 2. The methodology consists of four main modules: (1) Data Pre-processing Pipeline, (2) Dense Quantile Forecasting for Wind Power, (3) Dense Quantile Forecasting for Solar Power, and (4) Probabilistic Aggregation. The core idea is to model the probability density function (PDF) of wind and solar power through dense quantile forecasts, which are then aggregated to fit the distribution of total generation to obtain the target quantiles. LightGBM is employed as the primary model in the dense quantile forecasting for wind and solar power modules, as it is a popular tree-based model notably demonstrated in the M5 forecasting competition [32]. LightGBM is also recognized for its high training efficiency. According to our observations in offline experiments, the training time for a single quantile forecasting model is less than 30 seconds on the full dataset provided by HEFTCom2024 (i914900k CPU, 96GB memory). This choice helps improve the efficiency of pre-competition development and debugging processes, and is particularly effective in training multiple quantile regression models for fitting the CDF of wind and solar power. The details of each module are elaborated in Sections 3.2, 3.3, 3.4, and 3.5.
@@ 3.2 Data Pre-processing Pipeline
The data pre-processing pipeline is illustrated in Figure 3. We begin by processing the NWP data from the DWD and GFS models through spatial feature extraction, which reduces the dimensionality of the dense grid data. The extracted features include the mean, maximum, minimum, 25th percentile, and 75th percentile of the meteorological variables. Subsequently, we retain only NWP data with forecast lead time (FLT) within 48 hours, as the latest available FLT at the time of daily forecast submissions typically range from 23 to 47 hours. Including NWP data with FLT less than 23 hours can enhance forecast accuracy, while excluding those beyond 48 hours helps mitigate noise. Linear interpolation is employed to resample the NWP data to a half-hourly resolution. For the energy data, we reconstruct the actual available wind power based on the Bid Offer Acceptances (BOA) from the system operator. Subsequently, we perform outlier treatment on the power generation data from Hornsea 1 and solar power plants, which includes removing NaN values and filtering power values that exceed generation capacity. Once the energy and meteorological data are pre-processed, we align them to create the modeling dataset. In temporal feature engineering, we select time-shifted features for both past and future half-hour intervals. To determine the most relevant features, Random Forest (RF) [31] is used to rank feature importance. For wind power forecasting, 9 features are selected by combining the maximum, mean, and minimum wind speeds at 100 meters in the target region with lagged, leading, and current time steps. Similarly, for solar power forecasting, 13 features are selected by combining the maximum, mean, and minimum values of solar downward radiation and cloud cover in the target region, along with lagged, leading, and current time steps, as well as the hour-of-day. It is noteworthy that the final feature set is not strictly determined by the RF rankings. Manual selection is also performed based on train-test performance to determine the final features.
@@ 3.3 Wind Power Forecasting based on Stacked Sister Models
$r_{\text{DWD}.\tau,m}=-\frac{\partial L_{\tau}(y_{\text{w}.t},\hat{F}_{\text{DWD}.\tau}^{(m-1)}(\mathbf{x}_{\text{w}.t}))}{\partial\hat{F}_{\text{DWD}.\tau}^{(m-1)}(\mathbf{x}_{\text{w}.t})}\quad t=1,2,...,|\mathcal{W}_{\text{DWD}}|,$ (6)
@@
Building on this, LightGBM introduces improvements such as the histogram algorithm, leaf-wise best split algorithm, and parallel learning, which are not detailed here but can be found in the original paper [25]. The final wind power forecast is the stacked output of the two models, which can be represented as
@@
A special event that affects wind power forecasting is the planned outage of Hornsea 1 during the competition, which will limit the available capacity of wind power. This event has not been reflected in the offline dataset before. We design a quantile-wise truncation model to correct the wind power forecast results based on the available capacity of Hornsea 1, which is reported in the REMIT of ELEXON BSC website [33]. The corrected wind power forecast $\hat{y}_{\text{w}.\tau,t}^{{}^{\prime}}$ is obtained by

### v2

# A Hybrid Strategy for Probabilistic Forecasting and Trading of Aggregated Wind-Solar Power: Design and Analysis in HEFTCom2024

@@ Abstract
Obtaining accurate probabilistic energy forecasts and making effective decisions amid diverse uncertainties are routine challenges in future energy systems. This paper presents the winning solution of team GEB, which ranked 3rd in trading, 4th in forecasting, and 1st among student teams in the IEEE Hybrid Energy Forecasting and Trading Competition 2024 (HEFTCom2024). The solution provides accurate probabilistic forecasts for a wind-solar hybrid system, and achieves substantial trading revenue in the day-ahead electricity market. Key components include: (1) a stacking-based approach combining sister forecasts from various Numerical Weather Predictions (NWPs) to provide wind power forecasts, (2) an online solar post-processing model to address the distribution shift in the online test set caused by increased solar capacity, (3) a probabilistic aggregation method for accurate quantile forecasts of hybrid generation, and (4) a stochastic trading strategy to maximize expected trading revenue considering uncertainties in electricity prices. This paper also explores the potential of end-to-end learning to further enhance the trading revenue by shifting the distribution of forecast errors. Detailed case studies are provided to validate the effectiveness of these proposed methods. Code for all mentioned methods is available for reproduction and further research in both industry and academia.
@@ 1 Introduction
Accurately forecasting renewable energy production and enabling timely decision-making are crucial for the transition to resilient, low-carbon power systems (Petropoulos et al.,, 2022; Sweeney et al.,, 2020; Hong et al.,, 2020). To bridge the gap between academic research and industrial practice, the Hybrid Energy Forecasting and Trading Competition 2024 (HEFTCom2024) challenges participants to provide quantile forecasts for a hybrid power plant which consists of the Hornsea 1 offshore wind farm and solar capacity in East England, and to trade the energy in a day-ahead electricity market.
@@
Our performance can be attributed to two key factors. On one hand, the primary wind and solar power forecasts are provided by Gradient Boosting Decision Trees (GBDTs), which is a popular ensemble learning approach that has demonstrated strong performance in previous energy forecasting competitions, such as Global Energy Forecasting Competitions (GEFCom) 2017 and GEFCom2014 (Roach,, 2019; Smyl, Hua,, 2019; Nagy et al.,, 2016; Landry et al.,, 2016; Huang et al.,, 2016). We also incorporated several practical techniques including feature engineering with spatial and temporal meteorological information, hyperparameter optimization and ensemble learning. These techniques are widely recognized as effective strategies in forecasting practice (Landgraf,, 2019; Browell et al.,, 2020). On the other hand, building on these proven techniques, we designed series tailored approaches to address the specific challenges posed by HEFTCom2024. To clarify the motivation behind our approaches, the following introduce these challenges or characteristics of HEFTCom2024, along with our corresponding solutions.
@@
(1) Multiple NWP sources. To reduce the forecast variance, we developed a stacking-based ensemble method to combine sister forecasts trained on two different NWP sources provided by HEFTCom2024. Stacking is a widely used ensemble learning technique that improves forecasting performance by leveraging the complementary strengths of diverse base models (Ganaie et al.,, 2022). Stacking typically involves integrating models with different structures, such as combining Support Vector Regression (SVR) and Random Forest for day-ahead wind power forecasting (Ribeiro et al.,, 2022), or combining Long Short Term Memory (LSTM) and Informer networks for solar power forecasting (Cao et al.,, 2023). In contrast to heterogeneous ensembles, combining sister forecasts is also demonstrated effective in power load forecasting tasks (Nowotarski et al.,, 2016), where sister forecasts refer to predictions generated from a series of models that share similar architectures but are trained using different input sources or variable selections (Liu et al.,, 2015). In our solution, we stack sister forecasts trained on different NWP sources to exploit their spatial complementarity, as we observed discrepancies in grid point density and coverage area in the NWP data provided for the Hornsea 1 region, as shown in Section 2. Rather than directly merging multiple NWP sources into a single forecasting model, stacking sister models offers an additional advantage—it mitigates the risk of missing data from one NWP source during online testing. Since the two NWPs cover different spatial regions, embedding both into a single model would make the model vulnerable to data gaps if one NWP source becomes unavailable. In contrast, with a stacked ensemble, one sister model can still generate forecasts even when the other fails.
@@
(2) Solar distribution shift. The solar capacity in East England was observed from 2,609 MWp to 2,741 MWp as of February 19, 2024. Since the offline dataset only reflects capacities up to 2,609 MWp, the forecasting models solely trained on offline dataset become less effective at online test-set. Distribution shift is a typical challenge in forecasting practice. Common methods to address distribution shifts include oversampling, transfer learning, and post-processing. Oversampling replicates limited new samples to construct an augmented training set, which can alleviate overfitting to outdated data but may disrupt temporal continuity and lead to overfitting on the limited new samples or being overwhelmed by historical samples (Moniz et al.,, 2018; Zhang et al.,, 2022). Transfer learning aim to align the model outputs with the new data distribution through fine-tuning (Vrbančič et al.,, 2020) or domain adaptation (Kamath et al.,, 2019), but the structure of decision trees does not support typical transfer learning paradigms. Post-processing operates directly on model outputs, which is effective for linear shifts and does not require retraining the model. While they assume the existence of a known and simple shift pattern, such assumptions are valid in our case: the increase in solar capacity is both known and induces a near-linear expansion in power generation. We combined online learning with post-processing to dynamically fit the new distribution. Given the limited online samples, we adopted a polynomial post-processing model trained using LASSO-penalized quantile regression (Li & Zhu,, 2008) for calibration, which prevents overfitting of higher-order terms by penalizing the model complexity. While wind power forecasts may also be subject to distributional changes, such shifts tend to be less systematic and harder to characterize without prior knowledge. Applying nonlinear corrections to wind power forecasts under sparse online data conditions will introduce additional risks of overfitting. Consequently, we restrict the use of online post-processing to solar power forecasts in the final solution.
@@
(3) Requirement to provide quantile forecasts of total generation. Directly predicting the distribution of total generation would hinder targeted post-processing, as Hornsea 1 encountered planned outage that were not present in the offline dataset, and the distribution of solar power generation also exhibited a shift. Therefore, forecasting wind and solar power separately is necessary and also help reduces the capacity of the forecasting model. While summing the forecast results quantile by quantile offers simplicity, it theoretically sacrifices the accuracy. Discrete convolution can be used to combine multiple cumulative distribution functions (CDFs) to obtain the distribution of the sum (Laha & Rohatgi,, 2020). Copula-based discrete convolution methods are further researched to aggregate the uncertainties of wind power (Zhang et al.,, 2016) and load demand (Wang et al.,, 2018) in power system. Inspired by these theoretical frameworks and applications, we developed an aggregated probabilistic forecasting approach by first predicting the CDF of wind and solar power separately and then reconstructing the distribution of total generation via discrete convolution. The CDF is estimated by learning the dense quantiles. This approach allows for an accurate estimation of the overall generation distribution in a parameter-free manner. Furthermore, this approach benefits from powerful models like LightGBM (Ke et al.,, 2017) or CatBoost (Prokhorenkova et al.,, 2018), which directly support learning the pinball loss.
@@
(4) Requirement to trade energy in the day-ahead electricity market. The trading revenue is influenced by the uncertainties in day-ahead prices, imbalance prices, and power generation of the hybrid power plant. Due to limited information about the market and unknown market imbalances, price forecasting is challenging. However, statistically, the mean of the price spread follows a seasonal pattern throughout the day. To utilize this regularity, we developed a stochastic trading method based on uncertain price spread and deterministic power forecasts to maximize the expected trading revenue. We derived a revenue loss from the inaccuracies in both power forecasts and price spread forecasts to guide the training of the deterministic power forecasting model. This approach provides a simple and robust strategy that avoids the need for training price forecasting models. Additionally, we propose a further enhancement to improve the trading revenue through end-to-end learning (E2EL). E2EL refers to designing forecasting models with a focus on downstream decision-making tasks rather than solely minimizing statistical errors. E2EL is explored in related literature for unit commitment (Chen et al.,, 2021), economic dispatch (Han et al.,, 2021), and energy arbitrage (Sang et al.,, 2022) tasks. The key to achieving superior decision-making through E2EL lies in leveraging the asymmetric impact of forecast errors on downstream benefits, which is overlooked in accuracy-oriented forecasting. We also discovered this characteristic of error asymmetry in the market rules set by HEFTCom2024, and case studies demonstrate that E2EL can enhance trading revenue by adjusting the distribution preferences of forecast errors.
@@ 2.3 Provided Data
HEFTCom2024 provides two meteorological data sources for forecasting: Deutscher Wetterdienst (DWD) and the Global Forecast System (GFS). The grid coordinates of the historical meteorological data for Hornsea 1 and East England photovoltaic (PV) plants are shown in Figure 1. For Hornsea 1, 36 coordinates are provided from the DWD, and 9 coordinates are provided from the GFS. For the solar power plants, 20 coordinates are provided from both DWD and GFS sources. The data types include wind speed and wind direction at 100m and 10m heights, temperature, solar radiation, cloud cover, and other meteorological data, with hourly resolution. The data are corrected every six hours, resulting in multiple corrected data points for the same valid time. The historical power data for Hornsea 1 and East England PV plants are given at half-hour intervals (Browell et al.,, 2023). During the online testing phase, the wind power data for Hornsea 1 is typically updated with a one-week lag, while the solar power plant data is updated with a one-day lag. HEFTCom2024 also provides day-ahead and imbalance price data in Great Britain’s wholesale electricity market for the trading track. The aforementioned data span from September 20, 2020, to January 18, 2024. The evaluation period for the competition is from February 20, 2024, to May 19, 2024.
@@ 3.1 Overview
The overall framework of the developed methodology is illustrated in Figure 2. The methodology consists of four main modules: (1) Data Pre-processing Pipeline, (2) Dense Quantile Forecasting for Wind Power, (3) Dense Quantile Forecasting for Solar Power, and (4) Probabilistic Aggregation. The core idea is to model the probability density function (PDF) of wind and solar power through dense quantile forecasts, which are then aggregated to fit the distribution of total generation to obtain the target quantiles. LightGBM is employed as the primary model in the dense quantile forecasting for wind and solar power modules, as it is a popular tree-based model notably demonstrated in the M5 forecasting competition (Makridakis et al.,, 2016). LightGBM is also recognized for its high training efficiency. According to our observations in offline experiments, the training time for a single quantile forecasting model is less than 30 seconds on the full dataset provided by HEFTCom2024 (i914900k CPU, 96GB memory). This choice helps improve the efficiency of pre-competition development and debugging processes, and is particularly effective in training multiple quantile regression models for fitting the CDF of wind and solar power. The details of each module are elaborated in Sections 3.2, 3.3, 3.4, and 3.5.
@@ 3.2 Data Pre-processing Pipeline
The data pre-processing pipeline is illustrated in Figure 3. We begin by processing the NWP data from the DWD and GFS models through spatial feature extraction, which reduces the dimensionality of the dense grid data. The extracted features include the mean, maximum, minimum, 25th percentile, and 75th percentile of the meteorological variables. Subsequently, we retain only NWP data with forecast lead time (FLT) within 48 hours, as the latest available FLT at the time of daily forecast submissions typically range from 23 to 47 hours. Including NWP data with FLT less than 23 hours can enhance forecast accuracy, while excluding those beyond 48 hours helps mitigate noise. Linear interpolation is employed to resample the NWP data to a half-hourly resolution. For the energy data, we reconstruct the actual available wind power based on the Bid Offer Acceptances (BOA) from the system operator. Subsequently, we perform outlier treatment on the power generation data from Hornsea 1 and solar power plants, which includes removing NaN values and filtering power values that exceed generation capacity. Once the energy and meteorological data are pre-processed, we align them to create the modeling dataset. In temporal feature engineering, we select time-shifted features for both past and future half-hour intervals. To determine the most relevant features, Random Forest (RF) (Breiman,, 2001) is used to rank feature importance. For wind power forecasting, 9 features are selected by combining the maximum, mean, and minimum wind speeds at 100 meters in the target region with lagged, leading, and current time steps. Similarly, for solar power forecasting, 13 features are selected by combining the maximum, mean, and minimum values of solar downward radiation and cloud cover in the target region, along with lagged, leading, and current time steps, as well as the hour-of-day. It is noteworthy that the final feature set is not strictly determined by the RF rankings. Manual selection is also performed based on train-test performance to determine the final features.
@@ 3.3 Wind Power Forecasting based on Stacked Sister Models
$r_{\text{DWD}.\tau,m}=-\frac{\partial L_{\tau}(y_{\text{w}.t},\hat{F}_{\text{DWD}.\tau}^{(m-1)}(\mathbf{x}_{\text{w}.t}))}{\partial\hat{F}_{\text{DWD}.\tau}^{(m-1)}(\mathbf{x}_{\text{w}.t})}\quad t=1,2,.,|\mathcal{W}_{\text{DWD}}|,$ (6)
@@
Building on this, LightGBM introduces improvements such as the histogram algorithm, leaf-wise best split algorithm, and parallel learning, which are not detailed here but can be found in the original paper (Ke et al.,, 2017). The final wind power forecast is the stacked output of the two models, which can be represented as
@@
A special event that affects wind power forecasting is the planned outage of Hornsea 1 during the competition, which will limit the available capacity of wind power. This event has not been reflected in the offline dataset before. We design a quantile-wise truncation model to correct the wind power forecast results based on the available capacity of Hornsea 1, which is reported in the REMIT of ELEXON BSC website (REMIT,, 2024). The corrected wind power forecast $\hat{y}_{\text{w}.\tau,t}^{{}^{\prime}}$ is obtained by